Three charged objects are placed as follows: Object B (1.5mC), object R (2.0mC), and object L (unknown). The distance between B and R is the same as the distance between the R and L. If the force on R is zero, what is the charge on L?
1
Expert's answer
2019-04-15T10:07:53-0400
"F_{RB}=\frac{lq_Rq_B}{d^2}"

"F_{RL}=\frac{lq_Rq_L}{d^2}"

If the force on R is zero,


"F_{RB}=F_{RL}\implies\frac{lq_Rq_L}{d^2}=\frac{lq_Rq_B}{d^2}"

Thus,


"q_L=q_B=1.5mC"